# Bi-Hölder invariants in o-minimal structures

### v2

@@ Abstract.
As an application, we show that a complex analytic germ that is bi-$\alpha$-Hölder homeomorphic to the germ of a Euclidean space for some $\alpha$ sufficiently close to $1$ must be smooth. This provides a slightly stronger version of Sampaio’s smoothness theorem, in which the germs are assumed to be bi-$\alpha$-Hölder homeomorphic for every $\alpha\in(0,1)$.
@@ 1. Introduction
Motivated by these developments, we consider the following natural questions, ordered by decreasing difficulty, for definable germs $(X,0)$ and $(Y,0)$ in an o-minimal structure that are strongly bi-$\alpha$-Hölder equivalent:
@@
As an immediate application of (i) and (iii), we obtain a simple proof showing that if a complex analytic germ $(X,0)$ is bi-$\alpha$-Hölder equivalent to the germ of a $k$-plane for some $\alpha$ close to $1$, then $(X,0)$ must be smooth. This yields a slight strengthening of Sampaio’s smoothness theorem [Sampaio2025].
@@
  • For a definable set $\subset\mathbb{R}^{n}, X¯\overline{X}$ denotes its Euclidean closure. A closed definable germ at 0 means the germ at 0 of a closed definable set. If $X$ is a definable germ at $x_{0}\in\mathbb{R}^{n}$, the link of $X$ at $x_{0}$ is the set
@@ 2.1. Bi-$\alpha$-Hölder equivalence
The constant $L$ is referred to as the Hölder constant of $f$.
@@ 2.2. O-minimal structures
A set in $\mathcal{D}_{n}$ for some $n$ is called definable set. A map $f:A\to\mathbb{R}^{m}$ is called definable if its graph is a definable set.
@@
By similar arguments, one can construct a strong deformation retract from $p^{-1}([\varepsilon_{1},d])$ onto $p^{-1}([\varepsilon_{2},d])$ for every $d>\varepsilon_{2}$.
@@
Theorem 2.10 (Łojasiewicz inequality [Coste]).
@@ 3. Lipschitz normal embeddings
Let $(X,0)\subset\mathbb{R}^{n}$ and $(Y,0)\subset\mathbb{R}^{m}$ be definable set-germs. Suppose $(X,0)$ is LNE. There exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$ then $(Y,0)$ is also LNE.
@@
We will prove by contradiction that if $Y$ is not LNE then it cannot be bi-$\alpha$-Hölder to $(X,0)$ when $\alpha$ is sufficiently close to $1$.
@@
Let $\varphi:(X,0)\to(Y,0)$ be an bi-$\alpha$-Hölder homeomorphism. For points $x,y\in Y$ near 0, set $x^{\prime}:=\varphi^{-1}(x)$ and $y^{\prime}:=\varphi^{-1}(y)$. Let $\gamma_{x^{\prime}y^{\prime}}$ be a curve in $X$ joining $x^{\prime}$ and $y^{\prime}$ such that
${\rm length}(\gamma_{x^{\prime}y^{\prime}})\sim\tilde{d}(x^{\prime},y^{\prime}).$
@@
(2) $\tilde{d}(x,y)\lesssim{\rm length}(\varphi(\gamma_{x^{\prime}y^{\prime}}))\lesssim\big({\rm length}(\gamma_{x^{\prime}y^{\prime}})\big)^{\alpha}\sim\tilde{d}(x^{\prime},y^{\prime})^{\alpha}.$
@@
(3) $\tilde{d}(x^{\prime},y^{\prime})^{\alpha}\lesssim\|x^{\prime}-y^{\prime}\|^{\alpha}\lesssim\|x-y\|^{\alpha^{2}}.$
@@
When $\alpha$ is close enough to 11 we have $\beta<\alpha^{2}<1$, inequality (4) then contradicts (1). The proof is complete. ∎
@@ 4. Dimension of tangent cones
In this section, we prove that the dimension of the tangent cones of given definable germs must be the same if they are bi-$\alpha$-Hölder homeomorphic with when $\alpha$ sufficiently close to $1$. The proof is inspired by the arguments of Koike and Paunescu [Koike2009] in the bi-Lipschitz case.
@@
  • Let $x_{0}\in\overline{X}$. The dimension of the germ $(X,x_{0})$, denoted by $\dim(X,x_{0})$, is the dimension of $X\cap B(x_{0},r)$ with $r$ sufficiently small.
@@
It is known that if $X$ is definable then $C_{x_{0}}X$ is also definable, and moreover $\dim C_{x_{0}}(X)\leq\dim(X,x_{0})$.
@@
Since $\varphi$ is a bi-$\alpha$-Hölder homeomorphism, there exists a constant $K>0$ such that
@@
Let us denote $d_{S^{n-1}}(s,S_{A})$ the spherical distance to $S_{A}$. On $S^{n-1}$, in a small neighborhood $U$ of $S_{A}$, the functions ${\rm dist}(s,A)\sim d_{S^{n-1}}(s,S_{A})$. Therefore, there exist constants $\kappa_{1},\kappa_{2}>0$ and $\rho_{0}>0$ such that, for all $0<\rho\leq\rho_{0}$,
@@
We now estimate the measure of tubular neighborhoods of $S_{A}$ in $S^{n-1}$. The set $S_{A}$ is a definable subset of $S^{n-1}$ of dimension $\dim S_{A}=a-1$. We can stratify $S_{A}$ into finitely many $C^{1}$ embedded manifolds called strata; the strata of top dimension a1a-1 dominate the measure of small neighborhoods, while lower-dimensional strata contribute higher-order terms. Thus, there exist constants $c_{1},c_{2},\varepsilon_{0}>0$ such that, for all $0<\varepsilon\leq\varepsilon_{0}$,
@@
$|r_{s}^{-1}(y_{1})-r_{s}^{-1}(y_{2})|$ $=\big|H\big((1/s)\,H^{-1}(y_{1})\big)-H\big((1/s)\,H^{-1}(y_{2})\big)\big|$
$\leq\operatorname{Lip}(H)\cdot\big|(1/s)\,H^{-1}(y_{1})-(1/s)\,H^{-1}(y_{2})\big|$
$=\frac{\operatorname{Lip}(H)}{s}\,\big|H^{-1}(y_{1})-H^{-1}(y_{2})\big|$
$\leq\frac{\operatorname{Lip}(H)\operatorname{Lip}(H^{-1})}{s}\,|y_{1}-y_{2}|=\frac{C}{s}|y_{1}-y_{2}|.$
@@
Let $L$ be the Hölder constant of $h$. For $r>0$ small define
@@
Fix a sufficiently small $r$ and put $D:={\rm diam}(E)>0$ where ${\rm diam}(E)$ denotes the diameter of $E$. Let $\varepsilon>0$ be arbitrary. By outer regularity of Lebesgue measure and compactness of $\overline{E}$, there exists a finite family of closed axis-parallel cubes $\{Q_{i}\}_{i=1}^{N}$ of common side length $t$ with
@@
By Jung’s theorem, $h(Q_{i})$ is contained in a ball of radius $\sqrt{1/2}{\rm diam}(h(Q_{i}))$,
${\rm vol}_{n}\big(h(Q_{i})\big)\leq\omega_{n}\Big(\frac{{\rm diam}(h(Q_{i}))}{2}\Big)^{n}\leq\omega_{n}2^{-n/2}L^{n}n^{\frac{\alpha n}{2}}t^{\alpha n}.$
@@
(5) $\rho^{(\alpha-1)n}\lesssim r^{\frac{1}{\alpha}(\alpha-1)n}.$
@@
Since $X$ is definable, by Lemma 4.7, $X$ and $C_{0}(X)$ is ST-equivalent. That is, there are $d_{1},d_{2}>1$ and $C_{1},C_{2}>0$ such that $X\subset ST_{d_{1}}(C_{0}(X),C_{1})$ and $C_{0}(X)\subset ST_{d_{2}}(X,C_{2})$. Applying Lemma 4.6, we get
@@
for any $C_{1},C_{2}>0$ and any $1<d<d_{1}$ and $r$ sufficiently small. We now fix $d<d_{1}$. Choose $0<\alpha_{0}<1$ such that $d\alpha_{0}^{2}>1$. This implies that for all $\alpha_{0}\leq\alpha\leq 1$ we have $d\alpha^{2}>1$.
From Lemma 4.6, given $C_{3}>0$ and $d>1$, there are $C_{4},C_{5}>0$ such that
(8) $ST_{d/\alpha^{2}}(F,C_{4})\subset\varphi(ST_{d}(E,C_{3}))\subset ST_{d\alpha^{2}}(F,C_{5}).$
@@
(9) ${\rm vol}_{n}\bigg(\varphi(ST_{d\alpha^{2}}(E,C_{5}))\cap B(0,r)\bigg)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\bigg(ST_{d\alpha^{2}}(E,C_{5})\cap B(0,r)\bigg).$
Again by Lemma 4.6, there is a constant $C_{6}>0$ such that
$ST_{d}(F,C_{6})\subset\varphi(ST_{d\alpha^{2}}(E,C_{5})).$
It follows that
${\rm vol}_{n}(ST_{d}(F,C_{6})\cap B(0,r))\leq{\rm vol}_{n}(\varphi(ST_{d\alpha^{2}}(E,C_{5}))\cap B(0,r)).$
Then, by Lemma 4.10,
${\rm vol}_{n}(ST_{d\alpha^{2}}(E,C_{5})\cap B(0,r))\sim r^{(\alpha^{2}-1)d(n-a)}{\rm vol}_{n}(ST_{d}(E,C_{5})\cap B(0,r))$
@@
(10) $1\sim\frac{{\rm vol}_{n}(ST_{d}(F,C_{6})\cap B(0,r))}{{\rm vol}_{n}(ST_{d}(C_{0}(F),C_{6})\cap B(0,r))}\lesssim r^{K}\frac{{\rm vol}_{n}(ST_{d}(E,C_{5})\cap B(0,r))}{{\rm vol}_{n}(ST_{d}(C_{0}(F),C_{6})\cap B(0,r))}.$
@@
Set $E:=C_{0}(X)$ and $F:=\varphi(E)$. We will verify that the assumptions of Lemma 4.14 are satisfied. Since $\varphi(X)$ is definable, by Lemma 4.13, there exists $\alpha_{1}$ such that if $\alpha\in[\alpha_{1},1]$ then
$C_{0}(\varphi(X))=C_{0}(\varphi(C_{0}(X)))=C_{0}(F).$
Because $X$ is definable, it follows from the proof of Lemma 4.13 that $\varphi(X)\sim_{ST}\varphi(C_{0}(Y))=F$. Moreover, since $\varphi(X)$ is definable, Lemma 4.7 yields $\varphi(X)\sim_{ST}C_{0}(\varphi(X))=C_{0}(F)$. By the transitivity of the ST-equivalence relation, we obtain $F\sim_{ST}C_{0}(F)$. Applying Lemma 4.14, we deduce that
$\dim C_{0}(Y)=\dim C_{0}(\varphi(X))=\dim C_{0}(F)\leq\dim E=\dim C_{0}(X).$
@@
Write $h=(h_{1},\dots,h_{n})$ and $h^{-1}=(k_{1},\dots,k_{n})$. There exist $C,H>0$ such that for all $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$,
$\|h(x)-h(x^{\prime})\|\leq C\,\|x-x^{\prime}\|^{\alpha},\qquad\|h^{-1}(y)-h^{-1}(y^{\prime})\|\leq H\,\|y-y^{\prime}\|^{\alpha}.$
By the McShane–Whitney extension for Hölder functions (see for example [McShane1934]), each coordinate function $h_{i}:X\to\mathbb{R}$ admits an $\alpha$-Hölder extension $\widetilde{h}_{i}:\mathbb{R}^{n}\to\mathbb{R}$; and similarly each $k_{j}:Y\to\mathbb{R}$ extends to $\widetilde{k}_{j}:\mathbb{R}^{n}\to\mathbb{R}$. Set $\widetilde{h}=(\widetilde{h}_{1},\dots,\widetilde{h}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\widetilde{k}=(\widetilde{k}_{1},\dots,\widetilde{k}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then $\widetilde{h}|_{X}=h$ and $\widetilde{k}|_{Y}=h^{-1}$, and there are constants $C^{\prime},H^{\prime}>0$ such that
$\|\widetilde{h}(u)-\widetilde{h}(u^{\prime})\|\leq C^{\prime}\|u-u^{\prime}\|^{\alpha},\qquad\|\widetilde{k}(v)-\widetilde{k}(v^{\prime})\|\leq H^{\prime}\|v-v^{\prime}\|^{\alpha}\quad(u,u^{\prime},v,v^{\prime}\in\mathbb{R}^{n}).$
@@
$\leq\|u-u^{\prime}\|+H^{\prime}\,\bigl\|\,(v-v^{\prime})+(\widetilde{h}(u)-\widetilde{h}(u^{\prime}))\,\bigr\|^{\alpha}$
$\leq\|u-u^{\prime}\|+H^{\prime}\bigl(\|v-v^{\prime}\|+C^{\prime}\|u-u^{\prime}\|^{\alpha}\bigr)^{\alpha}$

### v3

@@ Abstract.
As an application, we give a simple proof that a complex analytic germ which is bi-$\alpha$-Hölder homeomorphic to the germ of a Euclidean space for some $\alpha$ sufficiently close to $1$ must be smooth. This provides a slightly stronger version of Sampaio’s smoothness theorem, in which the germs are assumed to be bi-$\alpha$-Hölder homeomorphic for every $\alpha\in(0,1)$.
@@ 1. Introduction
Motivated by these developments, we consider the following natural questions, listed in decreasing order of difficulty. Let $(X,0)$ and $(Y,0)$ be definable germs in an o-minimal structure and suppose they are strongly bi-$\alpha$-Hölder equivalent:
@@
As an immediate application of (i) and (iii), we obtain a simple proof showing that if a complex analytic germ $(X,0)$ is bi-$\alpha$-Hölder equivalent to the germ of a $k$-plane for some $\alpha$ close to $1$, then $(X,0)$ must be smooth. This yields a strengthening of Sampaio’s smoothness theorem [Sampaio2025].
@@
  • Let $X\subset\mathbb{R}^{n}$. We denote by ${\rm diam}(X)$ the diameter of $X$. For a point $x\in\mathbb{R}^{n}$, we denote by ${\rm dist}(x,X)$ the Euclidean distance from $x$ to $X$.
  • For a definable set $\subset\mathbb{R}^{n}, X¯\overline{X}$ denotes its Euclidean closure. A closed definable germ at $x_{0}$ means the germ at $x_{0}$ of a closed definable set. If $X$ is a definable germ at $x_{0}\in\mathbb{R}^{n}$, the link of $X$ at $x_{0}$ is the set
@@ Acknowledgements
We thank an anonymous referee for pointing out an error in the proof of Theorem 3.1 in an earlier version of this paper.
@@ 2.1. Bi-$\alpha$-Hölder equivalence
The constant $L$ is referred to as a Hölder constant of $f$.
@@ 2.2. O-minimal structures
A set in $\mathcal{D}_{n}$ for some $n$ is called a definable set. A map $f:A\to\mathbb{R}^{m}$ is called definable if its graph is definable.
@@
By similar arguments, one can construct a strong deformation retract from $p^{-1}([\varepsilon_{1},d])$ onto $p^{-1}([\varepsilon_{2},d])$ for every $d\geq\varepsilon_{2}$.
@@
Theorem 2.10 (Łojasiewicz inequality [Dries-Miller96]).
@@ 3. Lipschitz normal embeddings
Another important distance is the diameter distance, denoted by $d_{{\rm diam}}(x,y)$, defined as the infimum of the diameters of a curve in $X$ joining $x$ and $y$. It was proved in [Sampaio2025, Theorem 3.2] that, in o-minimal structures on $\mathbb{R}$ (not necessarily polynomially bounded), the distances $d_{\rm in}$ and $d_{{\rm diam}}$ are equivalent. As a consequence, we have $\tilde{d}$ and $d_{{\rm diam}}$ are equivalent.
@@
Let $(X,0)\subset\mathbb{R}^{n}$ and $(Y,0)\subset\mathbb{R}^{m}$ be definable germs. Suppose $(X,0)$ is LNE. There exists $\alpha_{0}\in(0,1)$ such that if $(X,0)$ and $(Y,0)$ are bi-$\alpha$-Hölder equivalent for some $\alpha\in[\alpha_{0},1]$ then $(Y,0)$ is also LNE.
@@
We will prove by contradiction that if $Y$ is not LNE then it cannot be bi-$\alpha$-Hölder homeomorphic to $(X,0)$ when $\alpha$ is sufficiently close to $1$.
@@
Let $\varphi:(X,0)\to(Y,0)$ be an bi-$\alpha$-Hölder homeomorphism. For points $x,y\in Y$ near 0, set $x^{\prime}:=\varphi^{-1}(x)$ and $y^{\prime}:=\varphi^{-1}(y)$. Since $\tilde{d}$ and $d_{{\rm diam}}$ are equivalent, we can choose a curve $\gamma_{x^{\prime}y^{\prime}}$ in $X$ joining $x^{\prime}$ and $y^{\prime}$ such that
${\rm diam}(\gamma_{x^{\prime}y^{\prime}})\sim\tilde{d}(x^{\prime},y^{\prime}).$
@@
(2) $\tilde{d}(x,y)\lesssim{\rm diam}(\varphi(\gamma_{x^{\prime}y^{\prime}}))\lesssim\big({\rm diam}(\gamma_{x^{\prime}y^{\prime}})\big)^{\alpha}\sim\tilde{d}(x^{\prime},y^{\prime})^{\alpha}.$
@@
(3) $\tilde{d}(x^{\prime},y^{\prime})^{\alpha}\sim\|x^{\prime}-y^{\prime}\|^{\alpha}\lesssim\|x-y\|^{\alpha^{2}}.$
@@
When $\alpha$ is close enough to $1$ we have $\beta<\alpha^{2}<1$, inequality (4) then contradicts (1). The proof is complete. ∎
@@ 4. Dimension of tangent cones
In this section, we prove that the dimension of the tangent cones of given definable germs must be the same if they are bi-$\alpha$-Hölder homeomorphic when $\alpha$ is sufficiently close to $1$. The proof is inspired by the arguments of Koike and Paunescu [Koike2009] in the bi-Lipschitz case.
@@
  • Let $x_{0}\in\overline{X}$. The dimension of the germ $(X,x_{0})$, denoted by $\dim(X,x_{0})$, is the dimension of $X\cap B(x_{0},r)$ where $r$ is sufficiently small.
@@
It is known that if $X$ is definable then $C_{x_{0}}(X)$ is also definable, and moreover $\dim C_{x_{0}}(X)\leq\dim(X,x_{0})$.
@@
Since $\varphi$ is a bi-$\alpha$-Hölder homeomorphism, there exists a constant $K\geq 1$ such that
@@
Let us denote $d_{S^{n-1}}(s,S_{A})$ the spherical distance to $S_{A}$. On $S^{n-1}$, in a small neighborhood $U$ of $S_{A}$, ${\rm dist}(s,A)\sim d_{S^{n-1}}(s,S_{A})$. Therefore, there exist constants $\kappa_{1},\kappa_{2}>0$ and $\rho_{0}>0$ such that, for all $0<\rho\leq\rho_{0}$,
@@
We now estimate the measure of tubular neighborhoods of $S_{A}$ in $S^{n-1}$. The set $S_{A}$ is a definable subset of $S^{n-1}$ of dimension a1a-1. We fix a finite stratification of $S_{A}$ into $C^{1}$ embedded submanifolds of $S^{n-1}$, called strata. Let $\Gamma$ be one of these strata. By Weyl’s tube formula (see, for example, [Weyl39]), for $\varepsilon>0$ sufficiently small,
${\rm vol}_{n-1}\!\left(\left\{s\in S^{n-1}\mid d_{S^{n-1}}(s,\Gamma)\leq\varepsilon\right\}\right)\sim\varepsilon^{\,n-1-\dim\Gamma}\,{\rm vol}_{\dim\Gamma}(\Gamma)\sim\varepsilon^{\,n-1-\dim\Gamma}.$
In particular, only the strata of maximal dimension a1a-1 contribute to the leading term of the volume of small tubular neighborhoods, while strata of lower dimension give rise to higher-order terms. Thus, there exist constants $c_{1},c_{2},\varepsilon_{0}>0$ such that, for all $0<\varepsilon\leq\varepsilon_{0}$,
@@
$\|r_{s}^{-1}(y_{1})-r_{s}^{-1}(y_{2})\|$ $=\big\|H\big((1/s)\,H^{-1}(y_{1})\big)-H\big((1/s)\,H^{-1}(y_{2})\big)\big\|$
$\leq\operatorname{Lip}(H)\cdot\big\|(1/s)\,H^{-1}(y_{1})-(1/s)\,H^{-1}(y_{2})\big\|$
$=\frac{\operatorname{Lip}(H)}{s}\,\big\|H^{-1}(y_{1})-H^{-1}(y_{2})\big\|$
$\leq\frac{\operatorname{Lip}(H)\operatorname{Lip}(H^{-1})}{s}\,\|y_{1}-y_{2}\|=\frac{C}{s}\|y_{1}-y_{2}\|.$
@@
Let $L\geq 1$ be a Hölder constant of $h$. For $r>0$ small define
@@
Fix a sufficiently small $r$ and put $D:={\rm diam}(E)>0$. Let $\varepsilon>0$ be arbitrary. By outer regularity of Lebesgue measure and compactness of $\overline{E}$, there exists a finite family of closed axis-parallel cubes $\{Q_{i}\}_{i=1}^{N}$ of common side length $t$ with
@@
By Jung’s theorem, $h(Q_{i})$ is contained in a ball of radius $2^{-1/2}{\rm diam}(h(Q_{i}))$,
${\rm vol}_{n}\big(h(Q_{i})\big)\leq\omega_{n}\Big(\frac{{\rm diam}(h(Q_{i}))}{2^{-1/2}}\Big)^{n}\leq\omega_{n}2^{-n/2}L^{n}n^{\frac{\alpha n}{2}}t^{\alpha n}.$
@@
(5) $\rho^{(\alpha-1)n}\leq(1/L)^{(\alpha-1)n}r^{\frac{1}{\alpha}(\alpha-1)n}.$
@@
Since $X$ is definable, by Lemma 4.7, $X$ and $C_{0}(X)$ is ST-equivalent. That is, there are $d_{1},d_{2}>1$ and $C_{1},C_{2}>0$ such that $X\subset ST_{d_{1}}(C_{0}(X),C_{1})$ and $C_{0}(X)\subset ST_{d_{2}}(X,C_{2})$ as germs at 0. Applying Lemma 4.6, we get
@@
for any $C_{1},C_{2}>0$ and any $1<d<d_{1}$ and $r$ sufficiently small. We now fix $1<d<d_{1}$. Choose $0<\alpha_{0}<1$ such that $d\alpha_{0}^{2}>1$. This implies that for all $\alpha_{0}\leq\alpha\leq 1$ we have $1<d\alpha^{2}<d$.
Fix a constant $C_{3}>0$, from Lemma 4.6, there is $C_{4}>0$ such that
$ST_{d}(F,C_{3})\subset\varphi(ST_{d\alpha^{2}}(E,C_{4})).$
It follows that
${\rm vol}_{n}(ST_{d}(F,C_{3})\cap B(0,r))\leq{\rm vol}_{n}(\varphi(ST_{d\alpha^{2}}(E,C_{4}))\cap B(0,r)).$
@@
(8) ${\rm vol}_{n}\bigg(\varphi(ST_{d\alpha^{2}}(E,C_{4}))\cap B(0,r)\bigg)\lesssim r^{(\frac{1}{\alpha}+1)(\alpha-1)n}{\rm vol}_{n}\bigg(ST_{d\alpha^{2}}(E,C_{4})\cap B(0,r)\bigg).$
Since $E$ is a definable cone at 0, by Lemma 4.10,
${\rm vol}_{n}(ST_{d\alpha^{2}}(E,C_{4})\cap B(0,r))\sim r^{(\alpha^{2}-1)d(n-a)}{\rm vol}_{n}(ST_{d}(E,C_{4})\cap B(0,r))$
@@
(9) $1\sim\frac{{\rm vol}_{n}(ST_{d}(F,C_{3})\cap B(0,r))}{{\rm vol}_{n}(ST_{d}(C_{0}(F),C_{3})\cap B(0,r))}\lesssim r^{K}\frac{{\rm vol}_{n}(ST_{d}(E,C_{4})\cap B(0,r))}{{\rm vol}_{n}(ST_{d}(C_{0}(F),C_{3})\cap B(0,r))}.$
@@
Set $E:=C_{0}(X)$ and $F:=\varphi(E)$. We will verify that the assumptions of Lemma 4.14 are satisfied. By Lemma 4.13, there exists $\alpha_{1}$ such that if $\alpha\in[\alpha_{1},1]$ then
(i) $\varphi(X)=Y\sim_{ST}\varphi(C_{0}(X))=F$
(ii) $C_{0}(Y)=C_{0}(F)$
By Lemma 4.7 we have (iii) $C_{0}(Y)\sim_{ST}Y$. By the transitivity of the ST-equivalence relation, we obtain $F\sim_{ST}C_{0}(F)$. Since $Y$ is definable, by (ii), $C_{0}(F)=C_{0}(Y)$ is also definable. Obviously, all assumptions in Lemma 4.14 are satisfied, so
$\dim C_{0}(Y)=\dim C_{0}(F)\leq\dim E=\dim C_{0}(X).$
@@
The following lemma is well known; see, for example, [Sampaio2025], where it appears in the proof of Theorem 4.1. For the reader’s convenience, we include a detailed proof here.
@@
Write $h=(h_{1},\dots,h_{n})$ and $h^{-1}=(k_{1},\dots,k_{n})$. There exist $C,L\geq 1$ such that for all $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$,
$\|h(x)-h(x^{\prime})\|\leq C\,\|x-x^{\prime}\|^{\alpha},\qquad\|h^{-1}(y)-h^{-1}(y^{\prime})\|\leq L\,\|y-y^{\prime}\|^{\alpha}.$
By the McShane–Whitney extension for Hölder functions (see for example [McShane1934]), each coordinate function $h_{i}:X\to\mathbb{R}$ admits an $\alpha$-Hölder extension $\widetilde{h}_{i}:\mathbb{R}^{n}\to\mathbb{R}$; and similarly each $k_{j}:Y\to\mathbb{R}$ extends to $\widetilde{k}_{j}:\mathbb{R}^{n}\to\mathbb{R}$. Set $\widetilde{h}=(\widetilde{h}_{1},\dots,\widetilde{h}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$ and $\widetilde{k}=(\widetilde{k}_{1},\dots,\widetilde{k}_{n}):\mathbb{R}^{n}\to\mathbb{R}^{n}$. Then $\widetilde{h}|_{X}=h$ and $\widetilde{k}|_{Y}=h^{-1}$, and there are constants $C^{\prime},L^{\prime}>\geq 1$ such that
$\|\widetilde{h}(u)-\widetilde{h}(u^{\prime})\|\leq C^{\prime}\|u-u^{\prime}\|^{\alpha},\qquad\|\widetilde{k}(v)-\widetilde{k}(v^{\prime})\|\leq L^{\prime}\|v-v^{\prime}\|^{\alpha}\quad(u,u^{\prime},v,v^{\prime}\in\mathbb{R}^{n}).$
@@
$\leq\|u-u^{\prime}\|+L^{\prime}\,\bigl\|\,(v-v^{\prime})+(\widetilde{h}(u)-\widetilde{h}(u^{\prime}))\,\bigr\|^{\alpha}$
$\leq\|u-u^{\prime}\|+L^{\prime}\bigl(\|v-v^{\prime}\|+C^{\prime}\|u-u^{\prime}\|^{\alpha}\bigr)^{\alpha}$